OpenStudy (anonymous):

What is the primefactorzation of 33? 1*33 3*11 3*13 3*3*3

OpenStudy (anonymous):

well which one equals 33?

OpenStudy (anonymous):

Which one do you think? It's not A because there are other numbers besides 1 and itself that can get into 33.

OpenStudy (anonymous):

there are 2, but then which isn't just 1* the number?

OpenStudy (holly00d1248):

both a and b im missing something arent i

OpenStudy (istim):

It's the second one. 1 is not a prime number.

OpenStudy (istim):

The others don't equal 33.

OpenStudy (holly00d1248):

O i got it never mind

OpenStudy (istim):

Therefore, the answer is 3*11.

OpenStudy (anonymous):

The two in the middle are sure not going to be it. Is there one on the top or the bottom that will equal 33?

OpenStudy (istim):

1 is not a prime number, thus it can't be a prime factorization of 33..

OpenStudy (istim):

The answer is B.

OpenStudy (holly00d1248):

i agree with @IsTim

OpenStudy (anonymous):

If you are just going to say the answer, then yes it is B.

OpenStudy (istim):

Guys, this is getting out of hand.. http://primes.utm.edu/lists/small/1000.txt THe only prime numbers relevant on the list, are 3, 11 and 13. 3* 11 = 33 3*13 = 39 (WRONG) 3*3*3 = 27 (Nope) and A doesn't even have all primes, therefore the answer must be B.